# On the Cutoff Scale Identification of FLRW Cosmology in Asymptotically Safe Gravity

### v1

@@ I Introduction
The effects of quantum gravity are crucial for establishing a consistent cosmological model spanning from the early to late universe. Developing a consistent and predictive theory of quantum gravity to achieve this is a significant challenge in theoretical physics. The perturbative approach in quantum gravity has been limited due to the negative mass dimension of the gravitational constant. However, in recent years, the asymptotically safe gravity has emerged as a promising scenario. The presence of a nontrivial fixed point renders the theory safe from UV divergences [1, 2]. The use of the functional renormalization group (FRG) within the context of asymptotically safe gravity has provided new avenues to explore phenomena in black hole physics and cosmology, extending from the UV scale to the IR scale [1, 3, 4, 5]. This approach enables the examination of consistent cosmological models across the entire UV to IR range.
Building on the framework of the FRG, the central element in the asymptotically safe gravity is the effective average action, $\Gamma_{k}[g_{\mu\nu}]$. This action is constructed to describe gravitational phenomena at a momentum scale $k$, accounting for the effects of quantum loops. To define the effective average action, a regulator term $R_{k}$ is introduced, which suppresses the contributions from momentum modes $p$ below $k$. As a result, quantum fluctuations below the FRG scale $k$ are excluded from the effective action, while modes with $p>k$ are fully integrated out. More explicitly the FRG equation is
$\frac{d\Gamma_{k}}{dt}=\frac{1}{2}{\rm Tr}\left[\Gamma_{k}^{(2)}+R_{k}\right]^{-1}\frac{dR_{k}}{dt},$ (1)
where $\Gamma_{k}^{(2)}$ is the second variation of the effective average action and $t=\log(k/k_{0})$ with $k_{0}$ being an arbitrary initial value.
@@
$S=\int d^{4}x\sqrt{-g}\frac{1}{16\pi G}(R-2\Lambda),$ (2)
where $G$ is the Newton constant and $\Lambda$ is the cosmological constant. The FRG ensures that the effective average action interpolates between the fundamental action, devoid of quantum corrections in the UV regime, and the quantum effective action as $k\rightarrow 0$. This interpolation makes the coupling constants scale dependent, and their evolution is governed by the FRG equation [1, 7, 6]. We show in the appendix A that the FRG equations yield scale-dependent expressions for these coupling constants:
@@
where $G_{0}$ and $\Lambda_{0}$ are the Newton coupling and cosmological constant at $k=0$. The dimensionless parameters $\omega,\omega_{1}$ etc. depend explicitly on the choice of regularization scheme, and we give the results for two schemes, the optimized and exponential cutoffs there. However the results are qualitatively similar, and both $G(k)$ and $\Lambda(k)$ may be written compactly to allow the $\Lambda_{0}=0$ solution by simply setting $\Lambda_{0}=0$ in (3). The solutions correspond to two distinct branches. The first branch, known as the Type IIa trajectory or the separatrix, assumes $\Lambda_{0}=0$ [8, 6, 9]. It extends to $k\rightarrow 0$, smoothly approaching the Gaussian fixed point in the IR regime. In contrast, the second branch, referred to as Type IIIa trajectories, corresponds to $\Lambda_{0}\neq 0$. This branch may not extend to $k\rightarrow 0$ due to the emergence of a singularity in the IR regime, limiting its validity to a finite momentum scale. However, as discussed in [10], the nonlocal extension of the Einstein-Hilbert truncation can potentially resolve the infrared singularity in the RG trajectories with a positive cosmological constant. It is crucial to note that the solutions for $\Lambda_{0}=0$ and $\Lambda_{0}\neq 0$ are derived separately, and we will see that these give distinct quantum corrections for the two branches. Nevertheless, the coefficients may be written uniformly like (3). We note that the quantum correction to the cosmological constant begins at the $k^{4}$ order for $\Lambda_{0}=0$, while it starts at the $k^{2}$-order for $\Lambda_{0}\neq 0$, which differs from previous reports in the literature [8, 11]. This distinction shows the need to consider quantum-improved cosmological solutions for the two cases separately.
@@
Quantum effects can be incorporated into cosmological studies at various levels. In this work, we incorporate quantum corrections into the Einstein equations by replacing the ordinary gravitational constant and cosmological constant with their scale-dependent counterparts. This approach allows us to explore the impact of quantum gravity on cosmology. A crucial step in this setup is the identification of the IR momentum scale $k$ with a suitable physical cutoff. Unlike black hole physics, where the cutoff scale can often be determined from physical mechanisms, cosmology lacks a direct method for such identification. Existing literature addresses this challenge by expressing $k$ in terms of physical quantities such as particle momenta or the spacetime curvature. Most natural and common choice in cosmology is to identify $k$ with the Hubble parameter, which is an almost unique physical scale. Various studies have explored Friedmann–Lemaître–Robertson–Walker (FLRW) cosmology choosing different cutoff scales [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. In this work, we aim to determine a suitable cutoff scale motivated by the classical solutions at late times. In general, different choices for the cutoff scale can lead to varying cosmic evolution scenarios. This raises a critical question: how can we identify which cutoff choices are more viable? To address this, we investigate the quantum-corrected late-time behaviour of the Hubble parameter and the scale factor. Since quantum corrections are expected to be small at late times, the quantum-corrected solutions should closely resemble the classical solutions with minor quantum improvements. This implies that the perturbations introduced by quantum corrections must not be so large as to significantly alter the form of the solutions at late times. Through this analysis, we can identify a way to assess the viability of different cutoff scale choices.
@@
We study the FLRW universe using the FRG-improved Einstein equations. This approach leads to a set of differential equations involving the Hubble parameter, energy density, the scale-dependent Newton’s gravitational constant, the cosmological constant, and a cutoff function. An important input for defining these equations is the requirement that the left-hand side of the Einstein equations is covariantly conserved, resulting in a modified continuity equation. This equation establishes a relationship between the matter density, the time-dependent gravitational constant, and the cosmological constant. Many cosmological studies have employed this modified continuity equation approach [16, 22, 22, 23, 20, 24]. It is important to note that this method differs from the consistency approach, which enforces the covariant conservation of both the energy-momentum tensor and the left-hand side of the Einstein equations [7, 12, 25]. In contrast, the modified continuity equation approach focuses on the conservation of the left-hand side of the Einstein equations only. Using this framework, we investigate the cosmological evolution for different cutoff choices across two trajectory branches: one with $\Lambda_{0}=0$ and the other with $\Lambda_{0}\neq 0$. For the $\Lambda_{0}=0$ branch, we consider cutoff identifications where $k$ is either the classical Hubble parameter or a function of the quantum Hubble parameter, and find that the choice of the classical Hubble parameter gives reasonable cosmology, and another choice of quantum Hubble parameter gets some constraint. For the $\Lambda_{0}\neq 0$ branch, the Hubble parameter becomes constant in late time, so it is not suitable to describe the time development. We find that the choice of the inverse cosmic time successful for $\Lambda_{0}=0$ is not suitable. Instead, we consider a time-dependent cutoff scale proportional to the inverse of the classical scale factor, $k\propto 1/a_{\mathrm{cl}}(t)$, which is suggested by the scalar curvature for compact space. We find that this choice ensure consistent solutions with the classical Hubble parameter and scale factor.
@@ II The Improved Field Equation
In this section, the FRG improvement is applied at the “equation level” by replacing the ordinary Newtonian gravitational constant and cosmological constant with the scale-dependent running coupling constants $G(k)$ and $\Lambda(k)$ in the Einstein equation. We take the viewpoint that these running coupling constants capture quantum effects at the leading order. The quantum improved Einstein equation can thus be written as
@@
In writing down this equation, $G(k)$ and $\Lambda(k)$ are replaced with $G(t)$ and $\Lambda(t)$ to facilitate the study of cosmology, where the infrared cutoff is identified as $k=k(t)$ with the cosmic time $t$.
@@
$\rho=-\frac{\partial_{k}\Lambda}{8\pi\partial_{k}G}.$ (13)
We see from the low energy expansions in Eq. (3), the leading term of $\partial_{k}G$ is negative (for $\omega>0$). So the sign of the energy density $\rho$ depends entirely on the sign of the coefficient of the leading quantum correction to the cosmological constant. For $\Lambda_{0}\neq 0$, it is negative, giving negative $\rho$. For $\Lambda_{0}=0$, it is positive (for $\nu>0$), resulting in positive $\rho$. The case of $\Lambda_{0}=0$ is often considered, and discussed in detail especially in Ref. [7]. However, in our case, there is another possibility, which we employ to identify the cutoff scale in both cases.
@@
When we consider the quantum effects, $G$ and $\Lambda$ depend on the cosmic time. This arises because of the quantum effects of gravity. It is then expected that they also contribute to the energy-momentum. It must be $G$ in front of the energy-momentum tensor and $\Lambda$ that give such contribution in the present formulation. This suggests that we should include the time derivative of $G(t)$ and $\Lambda(t)$ in Eq. (10) as the energy-momentum from the gravity, instead of making its both sides separately vanish. This modified equation has been employed in various studies of cosmology within the asymptotic safety framework [16, 22, 22, 23, 20, 24]. This is weaker condition than (i) and gives interesting cosmology for $\Lambda_{0}\neq 0$. In this paper, we take this condition.
@@
$\dot{H}=-\frac{1}{2}(3+3w)\left[H^{2}-\frac{1}{3}\Lambda(t)\right]=-\alpha\left(H^{2}-\frac{\Lambda}{3}\right)\,,$ (15)
where $\alpha\equiv 3(1+w)/2$. Furthermore, the energy density $\rho(t)$ can be expressed in terms of the solution for $H(t)$ and the cosmological constant $\Lambda(t)$:
$\rho=\frac{3}{8\pi G}\left(H^{2}-\frac{\Lambda}{3}\right).$ (16)
We shall solve the system of differential equations given by (15) and (16) for $H(t)$ and $\rho(t)$, under the condition that $H(t)\neq 0$ for the prescribed form of $G(t)$ and $\Lambda(t)$. It is important to note that to solve Eq. (15) for the Hubble parameter $H(t)$, only the form of the external function $\Lambda(t)$ is required. The running gravitational constant $G(t)$ only appears in Eq. (16) when solving for $\rho(t)$ in terms of $H(t)$ and $\Lambda(t)$. To determine $\Lambda(t)$, we have to make identification of the infrared momentum scale $k$ with the physical cutoff scale. Existing literature often expresses $k$ in terms of scales, such as particle momenta, field strengths, or the curvature of spacetime [7]. However, there is no satisfactory principles to directly correlate the FRG scale $k$ with the physical scale in cosmology, similar to the consistent thermodynamics for quantum improved black holes [26, 27]. For the FLRW universe, homogeneity and isotropy imply that the only physical sacle is the Hubble parameter which is determined by the cosmic time. Thus we have
@@
A natural choice of the cutoff identification is in terms of the Hubble parameter $H(t)$. We still have the possibility of using classical or quantum Hubble parameters. In the next section, we first consider FLRW cosmology for $\Lambda_{0}=0$ using the identification with the classical Hubble parameter, and then the quantum Hubble. In the latter case, we consider a general form of $k(H)=\xi G_{0}^{(\beta-1)/2}H^{\beta}$, where $\beta$ is constant.
@@ III.2 The identification with the classical Hubble parameter
Substituting the cosmological constant (III.2) into (15), and taking the ansatz for $H(t)$ as
@@
where $c$ is a new constant parameter. The appearance of one undetermined parameter $c$ is expected, as Eq. (15) is a first-order differential equation, which naturally provides one integration constant. We see the terms in the first bracket sums up to $1/(t-c)$, quite similar to the classical solution. However we cannot remove this parameter by absorbing it to the cosmic time $t$ because it changes the form of the cosmological constant term, in contrast to classical case. But in the absence of the quantum corrections ($\tilde{\nu}=\tilde{\nu}_{1}=0$), $c$ becomes identical with the classical parameter.
@@
The quantum parameter $\tilde{\nu}$, which defines the critical time, is always positive for both the optimized and exponential cutoffs: specifically, $\tilde{\nu}=\xi^{4}/8\pi$ for the former case, as defined in Eq. (90) and $\tilde{\nu}=\xi^{4}\zeta(3)/2\pi$ for the latter in Eq. (81). Therefore, this positive critical time scale indicates that the leading quantum correction has an attractive effect (AdS-phase) for which the expansion rate of the FLRW universe slows down compared with the speed of classical expansion, $H(t)<H_{\mathrm{cl}}(t)=1/\alpha t$, for $t<t_{c}$. Conversely, it becomes repulsive (dS-phase) and makes the expansion rate bigger than the classical expansion rate when $t>t_{c}$. The factor $(t_{c}/t-1)$ is not universal, but appears up to $t^{-3}$-order and the remaining terms are much smaller. So the critical time scale, which distinguishes the AdS and dS phases, is qualitatively correct.
@@
Note that, by setting $\tilde{\nu}=\tilde{\omega}=0$ and $c=0$, we recover the classical result of energy density in FLRW cosmology. Here we can have a further physical understanding for the existence of the critical time $t_{c}$. The leading-order quantum correction to the cosmological constant, $\Lambda\simeq\tilde{\nu}G_{0}/t^{4}>0$, introduces a repulsive force. However, the integration constant $c$ increases the energy density $\rho(t)$ which provides an attractive force. The entire effect is the competition of these two opposite contributions, with the critical time $t_{c}$ characterizing the balance point between them, as can be seen in Eq. (26).

### v2

# thanks: corresponding author

@@ I Introduction
The effects of quantum gravity are crucial for establishing a consistent cosmological model spanning from the early to late universe. Developing a consistent and predictive theory of quantum gravity to achieve this is a significant challenge in theoretical physics. The perturbative approach in quantum gravity has been limited due to the negative mass dimension of the Newton coupling. However, in recent years, the asymptotically safe gravity has emerged as a promising scenario. The presence of a nontrivial fixed point renders the theory safe from UV divergences [1, 2]. The use of the functional renormalization group (FRG) within the context of asymptotically safe gravity has provided new avenues to explore phenomena in black hole physics and cosmology, extending from the UV scale to the IR scale [1, 3, 4, 5]. This approach enables the examination of consistent cosmological models across the entire UV to IR range.
Building on the framework of the FRG, the central element in the asymptotically safe gravity is the effective average action, $\Gamma_{k}[g_{\mu\nu}]$. This action is constructed to describe gravitational phenomena at a momentum scale $k$, accounting for the effects of quantum loops. To define the effective average action, a regulator term $R_{k}$ is introduced, which suppresses the contributions from momentum modes $p$ below $k$. As a result, quantum fluctuations below the FRG scale $k$ are excluded from the effective action, while modes with $p>k$ are fully integrated out. More explicitly the FRG equation is [1, 5, 6]
$k\frac{d\Gamma_{k}}{dk}=\frac{1}{2}{\rm Tr}\left\{\left[\Gamma_{k}^{(2)}+R_{k}\right]^{-1}k\frac{dR_{k}}{dk}\right\},$ (1)
where $\Gamma_{k}^{(2)}$ is the second variation of the effective average action.
@@
$\Gamma_{k}=\frac{1}{16\pi G(k)}\int d^{4}x\sqrt{-g}(R(g)-2\Lambda(k)),$ (2)
where $R(g)$ is the Ricci scalar for the Riemannian metric $_{\mu\nu}, gg$ is $\det(g_{\mu\nu})$, $G(k)$ is the Newton coupling and $\Lambda(k)$ is the cosmological parameter. The FRG ensures that the effective average action interpolates between the fundamental action, devoid of quantum corrections in the UV regime, and the quantum effective action as $k\rightarrow 0$. This interpolation makes the coupling “constants” scale dependent, and their evolution is governed by the FRG equation [1, 8, 7]. For this reason, we call the Newton and cosmological constants Newton coupling and cosmological parameter. We show in the appendix A that the FRG equations yield scale-dependent expressions for these couplings:
@@
where $G_{0}$ and $\Lambda_{0}$ are the Newton coupling and cosmological parameter at $k=0$. The dimensionless parameters $\omega,\omega_{1}$ etc. depend explicitly on the choice of regularization scheme, and we give the results for two schemes, the optimized and exponential cutoffs there. However the results are qualitatively similar, and both $G(k)$ and $\Lambda(k)$ may be written compactly to allow the $\Lambda_{0}=0$ solution by simply setting $\Lambda_{0}=0$ in (3). The solutions correspond to two distinct branches. The first branch, known as the Type IIa trajectory or the separatrix, assumes $\Lambda_{0}=0$ [9, 7, 10]. It extends to $k\rightarrow 0$, smoothly approaching the Gaussian fixed point in the IR regime. In contrast, the second branch, referred to as Type IIIa trajectories, corresponds to $\Lambda_{0}\neq 0$. This branch may not extend to $k\rightarrow 0$ due to the emergence of a singularity in the IR regime, limiting its validity to a finite momentum scale. However, as discussed in [11], the nonlocal extension of the Einstein-Hilbert truncation can potentially resolve the infrared singularity in the RG trajectories with a positive cosmological parameter. It is crucial to note that the solutions for $\Lambda_{0}=0$ and $\Lambda_{0}\neq 0$ are derived separately, and we will see that these give distinct quantum corrections for the two branches. Nevertheless, the coefficients may be written uniformly like (3). We note that the quantum correction to the cosmological parameter begins at the $k^{4}$ order for $\Lambda_{0}=0$, while it starts at the $k^{2}$-order for $\Lambda_{0}\neq 0$, which differs from previous reports in the literature [9, 12]. This distinction shows the need to consider quantum-improved cosmological solutions for the two cases separately.
@@
Quantum effects can be incorporated into cosmological studies at various levels. In this work, we incorporate quantum corrections into the Einstein equations by replacing the ordinary Newton coupling and cosmological parameter with their scale-dependent counterparts. This approach allows us to explore the impact of quantum gravity on cosmology. A crucial step in this setup is the identification of the IR momentum scale $k$ with a suitable physical cutoff. Unlike black hole physics, where the cutoff scale can often be determined from physical mechanisms, cosmology lacks a direct method for such identification. Existing literature addresses this challenge by expressing $k$ in terms of physical quantities such as particle momenta or the spacetime curvature. Most natural and common choice in cosmology is to identify $k$ with the Hubble parameter, which is an almost unique physical scale. Various studies have explored Friedmann–Lemaître–Robertson–Walker (FLRW) cosmology choosing different cutoff scales [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In this work, we aim to determine a suitable cutoff scale motivated by the classical solutions at late times. In general, different choices for the cutoff scale can lead to varying cosmic evolution scenarios. This raises a critical question: how can we identify which cutoff choices are more viable? To address this, we investigate the quantum-corrected late-time behaviour of the Hubble parameter and the scale factor. Since quantum corrections are expected to be small at late times, the quantum-corrected solutions should closely resemble the classical solutions with minor quantum improvements. This implies that the perturbations introduced by quantum corrections must not be so large as to significantly alter the form of the solutions at late times. Through this analysis, we can identify a way to assess the viability of different cutoff scale choices.
@@
We study the FLRW universe using the FRG-improved Einstein equations. This approach leads to a set of differential equations involving the Hubble parameter, energy density, the scale-dependent Newton coupling, the cosmological parameter, and a cutoff function. An important input for defining these equations is the requirement that the left-hand side of the Einstein equations is covariantly conserved, resulting in a modified continuity equation. This equation establishes a relationship between the matter density, the time-dependent Newton coupling, and the cosmological parameter. Many cosmological studies have employed this modified continuity equation approach [17, 23, 24, 21, 25]. It is important to note that this method differs from the consistency approach, which enforces the covariant conservation of both the energy-momentum tensor and the left-hand side of the Einstein equations [8, 13, 26]. In contrast, the modified continuity equation approach focuses on the conservation of the left-hand side of the Einstein equations only. Using this framework, we investigate the cosmological evolution for different cutoff choices across two trajectory branches: one with $\Lambda_{0}=0$ and the other with $\Lambda_{0}\neq 0$. For the $\Lambda_{0}=0$ branch, we consider cutoff identifications where $k$ is either the classical Hubble parameter or a function of the quantum Hubble parameter, and find that the choice of the classical Hubble parameter gives reasonable cosmology, and another choice of quantum Hubble parameter gets some constraint. For the $\Lambda_{0}\neq 0$ branch, the Hubble parameter becomes constant in late time, so it is not suitable to describe the time development. We find that the choice of the inverse cosmic time successful for $\Lambda_{0}=0$ is not suitable. Instead, we consider a time-dependent cutoff scale proportional to the inverse of the classical scale factor, $k\propto 1/a_{\mathrm{cl}}(t)$, which is suggested by the scalar curvature for compact space. We find that this choice ensure consistent solutions with the classical Hubble parameter and scale factor.
@@ II The Improved Field Equation
In this section, the FRG improvement is applied at the “equation level” by replacing the ordinary Newton coupling and cosmological constant with the scale-dependent running coupling parameters $G(k)$ and $\Lambda(k)$ in the Einstein equation. We take the viewpoint that these running coupling parameters capture quantum effects at the leading order. The quantum improved Einstein equation can thus be written as
@@
In writing down this equation, $G(k)$ and $\Lambda(k)$ are replaced with $G(t)$ and $\Lambda(t)$ to facilitate the study of cosmology, where the infrared cutoff parameter becomes a function of the cosmic time $t$, so that $k=k(t)$.
@@
$\rho=-\frac{\partial_{k}\Lambda(k)}{8\pi\partial_{k}G(k)}.$ (13)
We see from the low energy expansions in Eq. (3), the leading term of $\partial_{k}G(k)$ is negative (for $\omega>0$). So the sign of the energy density $\rho$ depends entirely on the sign of the coefficient of the leading quantum correction to the cosmological parameter. For $\Lambda_{0}\neq 0$, it is negative, giving negative $\rho$. For $\Lambda_{0}=0$, it is positive (for $\nu>0$), resulting in positive $\rho$. The case of $\Lambda_{0}=0$ is often considered, and discussed in detail especially in Ref. [8]. However, in our case, there is another possibility, which we employ to identify the cutoff scale in both cases.
@@
When we consider the quantum effects, the Newton coupling and cosmological parameter depend on the cosmic time. This arises because of the quantum effects of gravity. It is then expected that they also contribute to the energy-momentum. It must be $G$ in front of the energy-momentum tensor and $\Lambda$ that give such contribution in the present formulation. This suggests that we should include the time derivative of $G(t)$ and $\Lambda(t)$ in Eq. (10) as the energy-momentum from the gravity, instead of making its both sides separately vanish. This modified equation has been employed in various studies of cosmology within the asymptotic safety framework [17, 23, 23, 24, 21, 25]. This is weaker condition than (i) and gives interesting cosmology for $\Lambda_{0}\neq 0$. In this paper, we take this condition.
@@
$\dot{H}=-\frac{1}{2}(3+3w)\left[H^{2}-\frac{1}{3}\Lambda(t)\right]=-\alpha\left(H^{2}-\frac{\Lambda(t)}{3}\right)\,$, (15)
where $\alpha\equiv 3(1+w)/2$. Furthermore, the energy density $\rho(t)$ can be expressed in terms of the solution for $H(t)$ and the cosmological parameter $\Lambda(t)$:
$\rho=\frac{3}{8\pi G(t)}\left(H^{2}-\frac{\Lambda(t)}{3}\right).$ (16)
We shall solve the system of differential equations given by (15) and (16) for $H(t)$ and $\rho(t)$, under the condition that $H(t)\neq 0$ for the prescribed form of $G(t)$ and $\Lambda(t)$. It is important to note that to solve Eq. (15) for the Hubble parameter $H(t)$, only the form of the external function $\Lambda(t)$ is required. The running Newton coupling $G(t)$ only appears in Eq. (16) when solving for $\rho(t)$ in terms of $H(t)$ and $\Lambda(t)$. To determine $\Lambda(t)$, we have to make identification of the infrared momentum scale $k$ with the physical cutoff scale. Existing literature often expresses $k$ in terms of scales, such as particle momenta, field strengths, or the curvature of spacetime [8]. However, there is no satisfactory principle to directly correlate the FRG scale $k$ with the physical scale in cosmology, similar to the consistent thermodynamics for quantum improved black holes [27, 28]. For the FLRW universe, homogeneity and isotropy imply that the only physical scale is the Hubble parameter which is determined by the cosmic time $t$. Thus we have
@@
A natural choice of the cutoff identification is in terms of the Hubble parameter $H(t)$. We still have the possibility of using classical or quantum Hubble parameters. In the next section, we first consider FLRW cosmology for $\Lambda_{0}=0$ using the identification with the classical Hubble parameter, and then the quantum Hubble. In the latter case, we consider a general form of $k(H)=\xi G_{0}^{(\beta-1)/2}H^{\beta}$, where $\xi$ and $\beta$ are constants.
@@ III.2 The identification with the classical Hubble parameter
Substituting the cosmological parameter (III.2) into (15), and taking the ansatz for $H(t)$ as
@@
where $c$ is a new constant parameter. The appearance of one undetermined parameter $c$ is expected, as Eq. (15) is a first-order differential equation, which naturally provides one integration constant. We see the terms in the first bracket sums up to 1/(tc)1/(t-c), quite similar to the classical solution. However we cannot remove this parameter by absorbing it to the cosmic time $t$ because it changes the form of the cosmological parameter term, in contrast to classical case. But in the absence of the quantum corrections ($\tilde{\nu}=\tilde{\nu}_{1}=0$), $c$ becomes identical with the classical parameter.
@@
The quantum parameter $\tilde{\nu}$, which defines the critical time, is always positive for both the optimized and exponential cutoffs: specifically, including the dimensionless parameter $\xi$, $\tilde{\nu}=\xi^{4}/8\pi$ for the former case, as defined in Eq. (91) and $\tilde{\nu}=\xi^{4}\zeta(3)/2\pi$ for the latter in Eq. (81). Therefore, this positive critical time scale indicates that the leading quantum correction has an attractive effect (AdS-phase) for which the expansion rate of the FLRW universe slows down compared with the speed of classical expansion, $H(t)<H_{\mathrm{cl}}(t)=1/\alpha t$, for $t<t_{c}$. Conversely, it becomes repulsive (dS-phase) and makes the expansion rate bigger than the classical expansion rate when $t>t_{c}$. The factor $(t_{c}/t-1)$ is not universal, but appears up to $t^{-3}$-order and the remaining terms are much smaller. So the critical time scale, which distinguishes the AdS and dS phases, is qualitatively correct.
@@
Note that, by setting $\tilde{\nu}=\tilde{\omega}=0$ and $c=0$, we recover the classical result of energy density in FLRW cosmology. Here we can have a further physical understanding for the existence of the critical time $t_{c}$. The leading-order quantum correction to the cosmological parameter, $\Lambda\simeq\tilde{\nu}G_{0}/t^{4}>0$, introduces a repulsive force. However, the integration constant $c$ increases the energy density $\rho(t)$ which provides an attractive force. The entire effect is the competition of these two opposite contributions, with the critical time $t_{c}$ characterizing the balance point between them, as can be seen in Eq. (26).